OpenStudy (anonymous):

What is the probability of drawing yellow marble followed by a red marble from a bag containing 12 yellow marbles, 14 red marbles, and 15 green marbles if the first marble is not replaced? 18/43 21/205 192/1,849

OpenStudy (anonymous):

\[\frac{\text{Number of Yellow Marbles}}{\text{Sum of Marbles}} \times \frac{\text{Number of Red Marbles}}{\text{Sum of Marbles - 1}} = ?\]

OpenStudy (anonymous):

so it would be 21/205?

OpenStudy (anonymous):

Yup :)